# Forget Vectors at Play: Universal Input Perturbations Driving Machine Unlearning in Image Classification

### v1

@@ 1 Introduction
To address (Q), we study MU from a fresh data-based viewpoint: forget vector, a universal input data perturbation designed to promote unlearning effectively; See the schematic overview in Fig. 1. Before developing the forget vector, we explore the rationale for how data-based perturbations complement current model-based MU approaches, as evidenced by these methods’ generalization to common data shifts, including Gaussian noise and adversarial perturbations Hendrycks & Dietterich (2019); Goodfellow et al. (2014). Our design for the forget vector is also related to recent input prompting techniques for vision models, known as visual prompting Bahng et al. (2022b); Chen et al. (2023); Oh et al. (2023) or model reprogramming Elsayed et al. (2018); Zhang et al. (2022); Chen (2024), used in transfer learning and model adaptation. These prompting methods learn input perturbations to enable a fixed model to perform well on new tasks, effectively guiding the model to execute tasks it wasn’t originally trained for. From this perspective, our research on the forget vector also explores whether it is possible to append a trainable “prompt” to the input to guide an already-trained neural network in unlearning specific data. We summarize our contributions below.
@@ 3 Preliminaries on MU and Problem Statement
In this work, we focus on the problem of MU for image classification. Let $\mathcal{D}=\{\mathbf{x}_{i},y_{i}\}_{i=1}^{N}$ represent a training set with $N$ examples, where $\mathbf{x}_{i}$ denotes the $i$th image data, and $y_{i}$ denotes its corresponding class label. We introduce a forget set $\mathcal{D_{\mathrm{f}}}\subseteq\mathcal{D}$, which specifies the training samples targeted for unlearning. Accordingly, the complement of $\mathcal{D_{\mathrm{f}}}$ is the retain set, i.e., $\mathcal{D_{\mathrm{r}}}=\mathcal{D}\setminus\mathcal{D_{\mathrm{f}}}$. The goal of MU is to efficiently and effectively eliminate the influence of $\mathcal{D_{\mathrm{f}}}$ on an already-trained model ${{\bm{\theta}}}$, so that the performance of the post-unlearning model closely approximates that of a model retrained from scratch on $\mathcal{D_{\mathrm{r}}}$ (i.e., excluding the impact of $\mathcal{D_{\mathrm{f}}}$ from scratch). Therefore, such a retraining method (referred to as Retrain) is typically considered as the glad standard of MU Thudi et al. (2022a); Jia et al. (2023). However, since Retrain is computationally intensive, most popular MU approaches instead address an unlearning optimization problem using the forget and retain sets to update the model parameters ${{\bm{\theta}}}$, starting from the originally pre-trained model, denoted by ${{\bm{\theta}}_{\mathrm{o}}}$. This yields the following optimization problem for MU:
@@
with the initialization ${\bm{\theta}}={{\bm{\theta}}_{\mathrm{o}}}$. In (2), $\ell_{\mathrm{MU}}$ represents an appropriate unlearning loss function that may depend on $\mathcal{D_{\mathrm{f}}}$ and/or $\mathcal{D_{\mathrm{r}}}$, as will be detailed when introducing specific unlearning methods. In the context of MU for image classification Golatkar et al. (2020); Fan et al. (2023), the specification of the forget set $\mathcal{D_{\mathrm{f}}}$ leads to two unlearning scenarios: class-wise forgetting, where $\mathcal{D_{\mathrm{f}}}$ consists of a subset focused on a specific image class targeted for unlearning, and random data forgetting, where $\mathcal{D}_{\mathrm{f}}$ is a randomly selected subset of images across all classes.
@@
In what follows, we introduce several representative MU approaches that serve as approximations to Retrain. (a) Fine-tuning (FT) Warnecke et al. (2021): This approach treats the MU problem as a continual learning task, defining the unlearning objective $\ell_{\mathrm{MU}}$ as a training objective that fine-tunes ${\bm{\theta}}$ over $\mathcal{D}_{\mathrm{r}}$ to induce catastrophic forgetting of $\mathcal{D}_{\mathrm{f}}$. (b) Random labeling (RL) Golatkar et al. (2020): This approach specifies the unlearning objective $\ell_{\mathrm{MU}}$ by assigning random labels or features to the data in $\mathcal{D}_{\mathrm{f}}$, thereby enforcing model forgetting. (c) Gradient ascent (GA) Thudi et al. (2022a): This approach employs the negative of the FT loss to reverse the training impact associated with the data in $\mathcal{D}_{\mathrm{f}}$. (d) Localization-informed unlearning Jia et al. (2023); Fan et al. (2023): This method identifies a subset of model weights critical to the unlearning task (e.g., through model sparsity Jia et al. (2023) or gradient saliency Fan et al. (2023)) and incorporates this weight localization as a prior to solve the unlearning problem in (2).
Given an unlearned model (denoted as ${{\bm{\theta}}_{\mathrm{u}}}$) after solving (2), unlearning performance is evaluated in two main areas: unlearning effectiveness, which measures whether the target data/information has been successfully removed, and utility retention, which assesses whether unlearning has preserved the model’s classification ability on unaffected data. Following the evaluation benchmark in Jia et al. (2023), unlearning effectiveness can be quantified by two metrics: unlearning accuracy (UA), defined as $1-$the model’s accuracy on $\mathcal{D_{\mathrm{f}}}$ (higher UA indicates better unlearning), and membership inference attack performance on $\mathcal{D_{\mathrm{f}}}$, termed MIA-Efficacy, where higher prediction accuracy on non-training samples indicates better unlearning. Utility retention is measured by retain accuracy (RA), reflecting the model’s accuracy on $\mathcal{D_{\mathrm{r}}}$, and testing accuracy (TA), which is the accuracy on the original test set.
@@
While previous MU approaches can be unified within the framework of (2) by varying the unlearning loss $\ell_{\mathrm{MU}}$ and weight localization priors, recent advancements in input data-based model adaptation, such as visual prompting Bahng et al. (2022a); Chen et al. (2023) and model reprogramming Chen (2024); Elsayed et al. (2018), suggest an alternative approach to MU. This strategy inspires us to design data-based prompting (implemented through universal input perturbations) to achieve unlearning without modifying the model itself. We refer to this input perturbation vector, designed specifically for MU, as the forget vector. To be more specific, let $\bm{\delta}$ represent the data-agnostic input perturbations to be designed. The problem of constructing a forget vector for MU can be formulated as
@@
where $\bm{\delta}$ is the perturbation variable, applied linearly to the forget samples as $\mathbf{x}^{\prime}\mathrel{\mathop{:}}=\mathbf{x}+\bm{\delta}$ for $\mathbf{x}\in\mathcal{D_{\mathrm{f}}}$, similar to visual prompting Bahng et al. (2022a) and adversarial examples Goodfellow et al. (2014). Unlike model-based MU methods, we will detail the unlearning objective function required for designing the forget vector in our later method sections.
@@ 4 Generalization of MU to Forget Data Shifts
Given an unlearned model ($\bm{\theta}_{\mathrm{u}}$) after solving (2), we examine two types of shifts in forget data: standard data corruptions used in the evaluation of OOD generalization Hendrycks et al. (2021); Hendrycks & Dietterich (2019) and worst-case perturbations generated by adversarial attacks Goodfellow et al. (2014); Madry et al. (2017).
(a) Data Corruptions. Following the OOD generalization evaluation approach in image classification Hendrycks et al. (2021); Hendrycks & Dietterich (2019), we introduce data corruptions from four main categories: noise, blur, weather, and digital. Each type of corruption includes five levels of severity, with higher levels representing increased noise intensity. Among these, we select zero-mean Gaussian noise (GN) and Elastic transformations (ET) as the primary corruption types to evaluate MU robustness against shifts in forget data. Our rationale is that Gaussian noise yields small pixel-wise perturbations (similar to adversarial perturbations introduced later) and Elastic transformations stretch or contract small image regions.
(b) Adversarial perturbations. An adversarial image is a benign image altered with carefully crafted, pixel-wise perturbations designed to mislead a classifier. In this work, we use the $\epsilon$-constrained $\ell_{\infty}$ norm-based $K$-step projected gradient descent (PGD) attack Goodfellow et al. (2014); Madry et al. (2017) to generate adversarial examples via iterative projected gradient updates. The parameter $\epsilon>0$ defines the radius of the $\ell_{\infty}$ norm of the perturbations, controlling their strength. And $K$ represents the number of PGD steps.
@@
Next, we apply the above data shift operations to the MU evaluation sets–namely, the forget, retain, and testing sets–and assess the unlearning performance of an unlearned model. Fig. 2 displays the performance of the gold standard unlearning method, Retrain, against Gaussian noise at test time with standard deviations of $0.08$ and $0.2$ Hendrycks & Dietterich (2019), and two types of Elastic transformations with parameters $\left(488,170.8,24.4\right)$ and $\left(488,19.52,48.8\right)$ regarding intensity, smoothing and offset for moderate and high-intensity distortions Hendrycks & Dietterich (2019), as well as a $7$-step PGD attack with perturbation strength $\epsilon=8/255$ Goodfellow et al. (2014). To ensure the feasibility of Retrain, we conduct the image classification task using ResNet-18 on the CIFAR-10 dataset.
@@ 5 Optimization for Forget Vectors
As motivated in Sec. 4, if we treat the forget vector as an external perturbation applied to the forget set, our goal becomes optimizing its design to maximize unlearning effectiveness without compromising the original model utility. To this end, we first propose an unlearning objective function, $\ell_{\mathrm{MU}}$, tailored for the forget vector design in (4), inspired by the untargeted C&W attack Carlini & Wagner (2017). We then introduce a novel paradigm called compositional unlearning, facilitated by forget vector arithmetic. These approaches center on optimizing the input perturbation $\bm{\delta}$ in (4) while keeping the model unchanged, i.e., identical to the pre-trained original model ${\bm{\theta}}={{\bm{\theta}}_{\mathrm{o}}}$.
@@
where $\lambda_{1}>0$ and $\lambda_{2}>0$ are the regularization parameters, and $\ell_{\mathrm{CE}}(\bm{\delta};{{\bm{\theta}}_{\mathrm{o}}},\mathcal{D_{\mathrm{r}}})$ denotes the CE loss of the model ${{\bm{\theta}}_{\mathrm{o}}}$ over the perturbed retain set $\{(\mathbf{x}+\bm{\delta},y)\}_{(\mathbf{x},y)\in\mathcal{D_{\mathrm{r}}}}$. Integrating (6) into (5), we can then apply stochastic gradient descent (SGD) Amari (1993) to optimize the forget vector variable $\bm{\delta}$.
@@
A forget vector defines an unlearning direction in the input space to guide the unlearning process. We explore whether a new unlearning direction can be efficiently constructed by interpolating from existing forget vectors, such as class-wise forget vectors obtained by solving (6) with $\mathcal{D}_{\mathrm{f}}$ defined as each class’s training set. This approach is analogous to the concept of task vectors in weight space for model editing Ilharco et al. (2022). However, to the best of our knowledge, input-based task vector arithmetic has not yet been explored in the literature. If forget vectors can be modified and combined using arithmetic operations, such as negation and addition, we can dynamically adjust a model’s unlearning behavior without re-solving the optimization problem (6) or any other model-based unlearning problem in (2). We refer to this new unlearning paradigm as compositional unlearning.
@@ 6.1 Experiment Setups
We focus on MU for image classification, using two datasets: CIFAR-10 Krizhevsky et al. (2009) and ImageNet-10, a 10-class subset of the original ImageNet (Deng et al., 2009), for ease of implementation of Retrain (exact unlearning) over ImageNet images as Huang et al. (2022); Tao et al. (2021). For these tasks, we use two well-trained image classifiers: ResNet-18 He et al. (2016) for CIFAR-10 and VGG-16 (Simonyan & Zisserman, 2014) for ImageNet-10 classification.
@@
In the context of MU for image classification, we consider two scenarios: class-wise forgetting and data-wise forgetting. In class-wise forgetting, training data from an image class are designated for unlearning, while in random data forgetting, a subset of all-class training points is randomly selected as the forget set, with a specified forget ratio of 10%. To demonstrate the effectiveness of our proposal, we consider 6 MU baseline methods, including ① FT Warnecke et al. (2021), ② RL Golatkar et al. (2020), ③ GA Thudi et al. (2022a), ④ NegGrad+ Kurmanji et al. (2024a), ⑤ SalUn Fan et al. (2023), and ⑥ SCRUB Kurmanji et al. (2024b).
@@
To solve the forget vector problem (4) with the proposed unlearning objective in (6), we set the retain loss regularization parameter $\lambda_{1}$ to 3 for CIFAR-10 and 5 for ImageNet-10 in class-wise forgetting, and to 1 for random data forgetting. The $\ell_{2}$-norm regularization parameter is set to $\lambda_{2}=1$. These hyperparameters are determined through a grid search over the range $[0,10]$. To optimize (4), we use stochastic gradient descent (SGD) Amari (1993) with a momentum factor of $0.9$ and an exponential learning rate scheduler, decaying at a rate of $0.9$ per iteration. Additionally, the batch size is set to $256$, with a maximum of 40 optimization iterations for both two datasets. To solve the compositional unlearning problem (7), we use a similar setup, setting both $\lambda_{1}$ and $\lambda_{2}$ to $1$.

### v2

@@ 1 Introduction
To address (Q), we study MU from a fresh data-based viewpoint: forget vector, a universal input data perturbation designed to promote unlearning effectively; See the schematic overview in Fig. 1. Before developing the forget vector, we explore the rationale for how data-based perturbations complement current model-based MU approaches, as evidenced by these methods’ generalization to common data shifts, including Gaussian noise and adversarial perturbations (Hendrycks & Dietterich, 2019; Goodfellow et al., 2014). Our design for the forget vector is also related to recent input prompting techniques for vision models, known as visual prompting (Bahng et al., 2022b; Chen et al., 2023; Oh et al., 2023) or model reprogramming (Elsayed et al., 2018; Zhang et al., 2022; Chen, 2024), used in transfer learning and model adaptation. These prompting methods learn input perturbations to enable a fixed model to perform well on new tasks, effectively guiding the model to execute tasks it wasn’t originally trained for. From this perspective, our research on the forget vector also explores whether it is possible to append a trainable “prompt” to the input to guide an already-trained neural network in unlearning specific data. We summarize our contributions below.
@@ 3 Preliminaries on MU and Problem Statement
In this work, we focus on the problem of MU for image classification. Let $\mathcal{D}=\{\mathbf{x}_{i},y_{i}\}_{i=1}^{N}$ represent a training set with $N$ examples, where $\mathbf{x}_{i}$ denotes the $i$th image data, and $y_{i}$ denotes its corresponding class label. We introduce a forget set $\mathcal{D_{\mathrm{f}}}\subseteq\mathcal{D}$, which specifies the training samples targeted for unlearning. Accordingly, the complement of $\mathcal{D_{\mathrm{f}}}$ is the retain set, i.e., $\mathcal{D_{\mathrm{r}}}=\mathcal{D}\setminus\mathcal{D_{\mathrm{f}}}$. The goal of MU is to efficiently and effectively eliminate the influence of $\mathcal{D_{\mathrm{f}}}$ on an already-trained model ${{\bm{\theta}}}$, so that the performance of the post-unlearning model closely approximates that of a model retrained from scratch on $\mathcal{D_{\mathrm{r}}}$ (i.e., excluding the impact of $\mathcal{D_{\mathrm{f}}}$ from scratch). Therefore, such a retraining method (referred to as Retrain) is typically considered as the glad standard of MU (Thudi et al., 2022a; Jia et al., 2023). However, since Retrain is computationally intensive, most popular MU approaches instead address an unlearning optimization problem using the forget and retain sets to update the model parameters ${{\bm{\theta}}}$, starting from the originally pre-trained model, denoted by ${{\bm{\theta}}_{\mathrm{o}}}$. This yields the following optimization problem for MU:
@@
with the initialization ${\bm{\theta}}={{\bm{\theta}}_{\mathrm{o}}}$. In (2), $\ell_{\mathrm{MU}}$ represents an appropriate unlearning loss function that may depend on $\mathcal{D_{\mathrm{f}}}$ and/or $\mathcal{D_{\mathrm{r}}}$, as will be detailed when introducing specific unlearning methods. In the context of MU for image classification (Golatkar et al., 2020; Fan et al., 2023), the specification of the forget set $\mathcal{D_{\mathrm{f}}}$ leads to two unlearning scenarios: class-wise forgetting, where $\mathcal{D_{\mathrm{f}}}$ consists of a subset focused on a specific image class targeted for unlearning, and random data forgetting, where $\mathcal{D}_{\mathrm{f}}$ is a randomly selected subset of images across all classes.
@@
In what follows, we introduce several representative MU approaches that serve as approximations to Retrain. (a) Fine-tuning (FT) (Warnecke et al., 2021): This approach treats the MU problem as a continual learning task, defining the unlearning objective $\ell_{\mathrm{MU}}$ as a training objective that fine-tunes ${\bm{\theta}}$ over $\mathcal{D}_{\mathrm{r}}$ to induce catastrophic forgetting of $\mathcal{D}_{\mathrm{f}}$. (b) Random labeling (RL) (Golatkar et al., 2020): This approach specifies the unlearning objective $\ell_{\mathrm{MU}}$ by assigning random labels or features to the data in $\mathcal{D}_{\mathrm{f}}$, thereby enforcing model forgetting. (c) Gradient ascent (GA) (Thudi et al., 2022a): This approach employs the negative of the FT loss to reverse the training impact associated with the data in $\mathcal{D}_{\mathrm{f}}$. (d) Localization-informed unlearning (Jia et al., 2023; Fan et al., 2023): This method identifies a subset of model weights critical to the unlearning task (e.g., through model sparsity (Jia et al., 2023) or gradient saliency (Fan et al., 2023)) and incorporates this weight localization as a prior to solve the unlearning problem in (2).
Given an unlearned model (denoted as ${{\bm{\theta}}_{\mathrm{u}}}$) after solving (2), unlearning performance is evaluated in two main areas: unlearning effectiveness, which measures whether the target data/information has been successfully removed, and utility retention, which assesses whether unlearning has preserved the model’s classification ability on unaffected data. Following the evaluation benchmark in (Jia et al., 2023), unlearning effectiveness can be quantified by two metrics: unlearning accuracy (UA), defined as $1-$the model’s accuracy on $\mathcal{D_{\mathrm{f}}}$ (higher UA indicates better unlearning), and membership inference attack performance on $\mathcal{D_{\mathrm{f}}}$, termed MIA-Efficacy, where higher prediction accuracy on non-training samples indicates better unlearning. Utility retention is measured by retain accuracy (RA), reflecting the model’s accuracy on $\mathcal{D_{\mathrm{r}}}$, and testing accuracy (TA), which is the accuracy on the original test set.
@@
While previous MU approaches can be unified within the framework of (2) by varying the unlearning loss $\ell_{\mathrm{MU}}$ and weight localization priors, recent advancements in input data-based model adaptation, such as visual prompting (Bahng et al., 2022a; Chen et al., 2023) and model reprogramming (Chen, 2024; Elsayed et al., 2018), suggest an alternative approach to MU. This strategy inspires us to design data-based prompting (implemented through universal input perturbations) to achieve unlearning without modifying the model itself. We refer to this input perturbation vector, designed specifically for MU, as the forget vector. To be more specific, let $\bm{\delta}$ represent the data-agnostic input perturbations to be designed. The problem of constructing a forget vector for MU can be formulated as
@@
where $\bm{\delta}$ is the perturbation variable, applied linearly to the forget samples as $\mathbf{x}^{\prime}\mathrel{\mathop{:}}=\mathbf{x}+\bm{\delta}$ for $\mathbf{x}\in\mathcal{D_{\mathrm{f}}}$, similar to visual prompting (Bahng et al., 2022a) and adversarial examples (Goodfellow et al., 2014). Unlike model-based MU methods, we will detail the unlearning objective function required for designing the forget vector in our later method sections.
@@ 4 Generalization of MU to Forget Data Shifts
Given an unlearned model ($\bm{\theta}_{\mathrm{u}}$) after solving (2), we examine two types of shifts in forget data: standard data corruptions used in the evaluation of OOD generalization (Hendrycks et al., 2021; Hendrycks & Dietterich, 2019) and worst-case perturbations generated by adversarial attacks (Goodfellow et al., 2014; Madry et al., 2017).
(a) Data Corruptions. Following the OOD generalization evaluation approach in image classification (Hendrycks et al., 2021; Hendrycks & Dietterich, 2019), we introduce data corruptions from four main categories: noise, blur, weather, and digital. Each type of corruption includes five levels of severity, with higher levels representing increased noise intensity. Among these, we select zero-mean Gaussian noise (GN) and Elastic transformations (ET) as the primary corruption types to evaluate MU robustness against shifts in forget data. Our rationale is that Gaussian noise yields small pixel-wise perturbations (similar to adversarial perturbations introduced later) and Elastic transformations stretch or contract small image regions.
(b) Adversarial perturbations. An adversarial image is a benign image altered with carefully crafted, pixel-wise perturbations designed to mislead a classifier. In this work, we use the $\epsilon$-constrained $\ell_{\infty}$ norm-based $K$-step projected gradient descent (PGD) attack (Goodfellow et al., 2014; Madry et al., 2017) to generate adversarial examples via iterative projected gradient updates. The parameter $\epsilon>0$ defines the radius of the $\ell_{\infty}$ norm of the perturbations, controlling their strength. And $K$ represents the number of PGD steps.
@@
Next, we apply the above data shift operations to the MU evaluation sets–namely, the forget, retain, and testing sets–and assess the unlearning performance of an unlearned model. Fig. 2 displays the performance of the gold standard unlearning method, Retrain, against Gaussian noise at test time with standard deviations of $0.08$ and $0.2$ (Hendrycks & Dietterich, 2019), and two types of Elastic transformations with parameters $\left(488,170.8,24.4\right)$ and $\left(488,19.52,48.8\right)$ regarding intensity, smoothing and offset for moderate and high-intensity distortions (Hendrycks & Dietterich, 2019), as well as a $7$-step PGD attack with perturbation strength $\epsilon=8/255$ (Goodfellow et al., 2014). To ensure the feasibility of Retrain, we conduct the image classification task using ResNet-18 on the CIFAR-10 dataset.
@@ 5 Optimization for Forget Vectors
As motivated in Sec. 4, if we treat the forget vector as an external perturbation applied to the forget set, our goal becomes optimizing its design to maximize unlearning effectiveness without compromising the original model utility. To this end, we first propose an unlearning objective function, $\ell_{\mathrm{MU}}$, tailored for the forget vector design in (4), inspired by the untargeted C&W attack (Carlini & Wagner, 2017). We then introduce a novel paradigm called compositional unlearning, facilitated by forget vector arithmetic. These approaches center on optimizing the input perturbation $\bm{\delta}$ in (4) while keeping the model unchanged, i.e., identical to the pre-trained original model ${\bm{\theta}}={{\bm{\theta}}_{\mathrm{o}}}$.
@@
where $\lambda_{1}>0$ and $\lambda_{2}>0$ are the regularization parameters, and $\ell_{\mathrm{CE}}(\bm{\delta};{{\bm{\theta}}_{\mathrm{o}}},\mathcal{D_{\mathrm{r}}})$ denotes the CE loss of the model ${{\bm{\theta}}_{\mathrm{o}}}$ over the perturbed retain set $\{(\mathbf{x}+\bm{\delta},y)\}_{(\mathbf{x},y)\in\mathcal{D_{\mathrm{r}}}}$. Integrating (6) into (5), we can then apply stochastic gradient descent (SGD) (Amari, 1993) to optimize the forget vector variable $\bm{\delta}$.
@@
A forget vector defines an unlearning direction in the input space to guide the unlearning process. We explore whether a new unlearning direction can be efficiently constructed by interpolating from existing forget vectors, such as class-wise forget vectors obtained by solving (6) with $\mathcal{D}_{\mathrm{f}}$ defined as each class’s training set. This approach is analogous to the concept of task vectors in weight space for model editing (Ilharco et al., 2022). However, to the best of our knowledge, input-based task vector arithmetic has not yet been explored in the literature. If forget vectors can be modified and combined using arithmetic operations, such as negation and addition, we can dynamically adjust a model’s unlearning behavior without re-solving the optimization problem (6) or any other model-based unlearning problem in (2). We refer to this new unlearning paradigm as compositional unlearning.
@@ 6.1 Experiment Setups
We focus on MU for image classification, using two datasets: CIFAR-10 (Krizhevsky et al., 2009) and ImageNet-10, a 10-class subset of the original ImageNet (Deng et al., 2009), for ease of implementation of Retrain (exact unlearning) over ImageNet images as (Huang et al., 2022; Tao et al., 2021). For these tasks, we use two well-trained image classifiers: ResNet-18 (He et al., 2016) for CIFAR-10 and VGG-16 (Simonyan & Zisserman, 2014) for ImageNet-10 classification.
@@
In the context of MU for image classification, we consider two scenarios: class-wise forgetting and data-wise forgetting. In class-wise forgetting, training data from an image class are designated for unlearning, while in random data forgetting, a subset of all-class training points is randomly selected as the forget set, with a specified forget ratio of 10%. To demonstrate the effectiveness of our proposal, we consider 6 MU baseline methods, including ① FT (Warnecke et al., 2021), ② RL (Golatkar et al., 2020), ③ GA (Thudi et al., 2022a), ④ NegGrad+ (Kurmanji et al., 2024a), ⑤ SalUn (Fan et al., 2023), and ⑥ SCRUB (Kurmanji et al., 2024b).
@@
To solve the forget vector problem (4) with the proposed unlearning objective in (6), we set the retain loss regularization parameter $\lambda_{1}$ to 3 for CIFAR-10 and 5 for ImageNet-10 in class-wise forgetting, and to 1 for random data forgetting. The $\ell_{2}$-norm regularization parameter is set to $\lambda_{2}=1$. These hyperparameters are determined through a grid search over the range $[0,10]$. To optimize (4), we use stochastic gradient descent (SGD) (Amari, 1993) with a momentum factor of $0.9$ and an exponential learning rate scheduler, decaying at a rate of $0.9$ per iteration. Additionally, the batch size is set to $256$, with a maximum of 40 optimization iterations for both two datasets. To solve the compositional unlearning problem (7), we use a similar setup, setting both $\lambda_{1}$ and $\lambda_{2}$ to $1$.